# Transformer Block Coupling and its Correlation with Generalization in LLMs

### v3

@@ Abstract
Large Language Models (LLMs) have made significant strides in natural language processing, and a precise understanding of the internal mechanisms driving their success is essential. In this work, we trace the trajectories of individual tokens as they pass through transformer blocks, and linearize the system along these trajectories through their Jacobian matrices. By examining the relationships between these Jacobians, we uncover the transformer block coupling phenomenon in a multitude of LLMs, characterized by the coupling of their top singular vectors across tokens and depth. Our findings reveal that coupling positively correlates with model performance, and that this relationship is stronger than with other hyperparameters, namely parameter budget, model depth, and embedding dimension. We further investigate the emergence of these properties through training, noting the development of coupling, as well as an increase in linearity and layer-wise exponential growth in the token trajectories. These collective insights provide a novel perspective on the interactions between token embeddings, and prompt further approaches to study training and generalization in LLMs.
@@ 1 Introduction
Transformers (Vaswani et al., 2017) can be represented as discrete, nonlinear, coupled dynamical systems, operating in high dimensions (Greff et al., 2016; Papyan et al., 2017; Haber & Ruthotto, 2017; Ee, 2017; Ebski et al., 2018; Chen et al., 2018; Bai et al., 2019; Rothauge et al., 2019; Gai & Zhang, 2021; Li & Papyan, 2023). Viewing the skip connections as enabling a discrete time step, we represent the hidden representations as dynamically evolving through the layers of the network. The term nonlinear refers to the nonlinear transformations introduced by activation functions, and coupled refers to the interdependent token trajectories that interact through the MLP and self-attention blocks.
In our work, we investigate whether there are identifiable structural characteristics across 38+ pretrained LLMs, measure their emergence with training, and analyze their relationship with generalization performance. During inference, as token embeddings pass through the network, we linearize the effect of transformer blocks on the token embeddings throughout the depth of the LLM. To this end, we compute the Jacobians of distinct connections between layers or tokens, derive their singular value decompositions (SVDs), and compare the resulting singular vectors. This approach measures the degree of coupling between singular vectors to capture the operational similarity of blocks as they act on tokens. This perspective raises several questions:
@@
Figure 1: Transformer Coupling Measurements. (a) The plot illustrates the correlation between average coupling (negative mis-coupling) and benchmark scores across base LLMs (not fine-tuned), showing that higher alignment corresponds to improved performance, with a regression fit yielding an $R^{2}$ value of 0.75 with a significant p-value of $1.56\times 10^{-6}$. (b) The mean normalized coupling (Section 3.1) is plotted as a function of training checkpoints for Pythia 12B and 6.9B (Biderman et al., 2023), measured at steps ${128,256,512,1k,2k,\ldots,128k,143k}$. (c-e) Adjacency plots illustrate the mean coupling scores between pairs of layers. Each node represents a layer, and edge weight and opacity indicate the strength of depth-wise normalized coupling. Visualizations are provided for checkpoints 1, 4k, and 143k of Pythia 12B.
@@ 1.1 Contributions
We investigate the motivating questions across several openly-available LLMs, most having over 1B parameters, trained by 6 independent organizations, with varying training methods and data (Appendix A.1). Through our experiments, we identify consistent patterns that define the transformer block coupling phenomenon.
1. Coupling. The singular vectors of the Jacobians of the transformer blocks are coupled across depth (Figures 3, LABEL:fig:paris_all_8to16_appendix) and tokens (Figures 4, LABEL:fig:coupling_fixedin, LABEL:fig:coupling_fixedout, LABEL:fig:coupling_inoutsame_appendix, LABEL:fig:coupling_fixed_input_appendix, LABEL:fig:coupling_fixed_output_appendix) in several open source LLMs (Table (1)). Furthermore, coupling across Jacobians emerges with training (Figures 1b, 3, 4, LABEL:fig:coupling_fixedin, LABEL:fig:coupling_fixedout), and the coupling strength becomes more pronounced between adjacent layers with training, indicating a layer-wise locality in the interactions (Figures 1(c-e)).
2. Generalization. The strength of coupling is correlated with benchmark performance on the HuggingFace Open LLM Leaderboard (Beeching et al., 2023) (Figures 1a, LABEL:fig:alignment_vary_k). Additionally, coupling is more strongly correlated with generalization than parameter budget, model depth, and token embedding dimension (Figure LABEL:fig:correlation_hyperparams).
3. Regularity. Linearity in hidden trajectories emerges with training (Figures LABEL:figs:lss-training, LABEL:figs:pca, LABEL:fig:lss-all, 17, 24, 25), aligning with behaviour previously observed in ResNets Li & Papyan (2023). Exponential growth occurs in contiguous token representations as a function of depth (Figures LABEL:fig:expo-all, 23, 24, LABEL:figs:expo-training), starkly contrasting linear growth previously observed.
We provide a new perspective on token embedding interactions within LLMs by examining layers of a transformer through their Jacobian matrices. Our results display the effect of training on transformer blocks, and suggest potential approaches for promoting generalization in LLMs.
@@ 2 Background on Large Language Models
We describe LLMs as a deep composition of functions that iteratively transform token embeddings. In the input layer, $l=0$, the text prompts are tokenized and combined with the positional encodings to create an initial high-dimensional embedding, denoted by $x_{i}^{0}\in\mathbb{R}^{d_{\text{model}}}$ for the $i^{\text{th}}$ token. When these embeddings are stacked, they form a matrix:
@@ 3.1 Coupling of Singular Vectors of Jacobians
$J_{t_{1}t_{2}}^{l}=U_{l}S_{l}V_{l}^{\top}$ and $J_{t_{1}^{\prime}t_{2}^{\prime}}^{l^{\prime}}=U_{l^{\prime}}S_{l^{\prime}}V_{l^{\prime}}^{\top},$ (9)
@@
$A_{t_{1}t_{2}t_{1}^{\prime}t_{2}^{\prime}}^{ll^{\prime}}$ :-$\coloneq U_{l^{\prime}}^{\top}J_{t_{1}t_{2}}^{l}V_{l^{\prime}}:$ (10)
$=U_{l^{\prime}}^{\top}U_{l}S_{l}V_{l}^{\top}V_{l^{\prime}}$ (11)
for $l,l^{\prime}\in\{1,\ldots,L\}$ and $t_{1},t_{2},t_{1}^{\prime},t_{2}^{\prime}\in\{1,\ldots,n\}$. If the singular vectors of distinct Jacobians are strongly aligned, then
$U_{l^{\prime}}^{\top}U_{l}\approx I\approx V_{l}^{\top}V_{l^{\prime}},$ (12)
implying that the coupling matrix $A$ should be strongly diagonal. Explicitly, we quantify the mis-coupling of $A$ using
$m(A)=\lVert A-\text{Diag}(A)\rVert_{F},$ (13)
where $\text{Diag}(A)$ is the matrix $A$ with all non-diagonal entries replaced by zero and $\lVert\cdot\rVert_{F}$ denotes the Frobenius norm. For normalized comparison between models, we normalize by $A$;
$\tilde{m}(A)=\frac{\lVert A-\text{Diag}(A)\rVert_{F}}{\lVert A\rVert_{F}}.$ (14)
@@ 3.3 Layer-Wise Exponential Growth
This measurement is motivated by the discussion in Section 6.1 the parametrization discussed in Appendix A.5, as well as empirical evidence in Figure LABEL:figs:norm-expo-traininga, and serves as a method to test the validity of the linearization presented in Equation 19.
@@ 3.4 Vision Transformer Training
For further investigation of coupling in transformers, we train a series of Vision Transformers (ViTs) following DEiT training (Touvron et al., 2021). We train 64 ViTs on CIFAR10 (Krizhevsky, 2009) with varied weight decay and stochastic depth rate for a fixed architecture of embedding size 192, depth 12, and 3 attention heads. Please see Appendix A.7 for further details.
Figure 3: Transformer Block Coupling across Depth. The figure shows Jacobian coupling across transformer blocks 9 to 16, using the prompt ”What is the capital of France? The capital is” to trace the final token’s trajectory. In trained models (bottom row), the diagonal pattern with minimal off-diagonal values indicates alignment of Jacobians, where top singular vectors of $J^{l^{\prime}}$ diagonalize $J^{l}$. Untrained models (top row) lack this alignment. Further details are in the Appendix. A.8 (Figure LABEL:fig:paris_all_8to16_appendix). Best viewed in color.
@@ 4.1 Suite of Large Language Models
Our study evaluates a total of 38 LLMs (24 base LLMs and 14 fine-tuned, see Appendix A.1) that were independently trained by various individuals and organizations. These models, provided through HuggingFace (Wolf et al., 2020), vary in terms of parameter budgets, number of layers, hidden dimensions, and training tokens. Moreover, we analyze the dynamics of each measurement throughout training by deploying the Pythia Scaling Suite (Biderman et al., 2023). A summary of the models under consideration is presented in Table 1 of Appendix A.1 and further details in Appendix A.6.
@@ 4.2 Prompt Data
We evaluate these LLMs using prompts of varying length, ambiguity, and context, sourced from the test set of ARC (Clark et al., 2018), GSM8K (Cobbe et al., 2021), HellaSwag (Zellers et al., 2019), MMLU (Hendrycks et al., 2021), Truthful QA (Lin et al., 2022), and Winogrande (Sakaguchi et al., 2019). This data sets the performance benchmarks on the HuggingFace Open LLM Leaderboard (Beeching et al., 2023) and provide a representative evaluation of performance on many language tasks.
@@ 5.1 Coupling of Jacobians Across Depth
In trained LLMs, we observe coupling of the top singular vectors of the Jacobians across depth (Figure 3 bottom row), evident in the low non-diagonal values with a visible diagonal present in the matrix subplots. This is consistently observed across various LLMs considered. On the other hand, in untrained models (Figure 3 top row), there is no coupling of Jacobians across different depths. There is coupling along the diagonal, however, because each Jacobian is trivially diagonalized by its own singular vectors. This, in addition to Figure 1, suggests that coupling across depth emerges through training.
@@ 5.2 Coupling of Jacobians Across Tokens
We analyze the coupling of singular vectors of Jacobians across tokens. For input and output tokens that are the same ($J_{l}^{tt}$ and $J_{l^{\prime}}^{t^{\prime}t^{\prime}}$, Figure 4), we observe strong coupling, indicating that a token’s interactions along its trajectory are coupled with others. For context tokens, coupling is examined by fixing the input token ($J_{l}^{t_{1}t_{2}}$ and $J_{l^{\prime}}^{t_{1}t_{2}^{\prime}}$, Figure LABEL:fig:coupling_fixedin) or the output token ($J_{l}^{t_{1}t_{2}}$ and $J_{l^{\prime}}^{t_{1}^{\prime}t_{2}}$, Figure LABEL:fig:coupling_fixedout). While coupling exists, it is less consistent across pairs. Untrained models (Figure 4 top row) show no such coupling.
Figure 4: Transformer Block Coupling across Tokens (Same input and output tokens). The figure shows Jacobian coupling for the same input and output token across tokens, visualized using the absolute values of $A_{ll^{\prime}}^{ttt^{\prime}t^{\prime}}$ (with fixed layers $l,l^{\prime}$). In trained models (bottom row), the strong diagonal and small off-diagonal values indicate coupling, while no such coupling is present at initialization (top row). Additional details are in Appendix A.8 (Figure LABEL:fig:coupling_inoutsame_appendix).
@@ 5.3 Emergence of Coupling with Training
Coupling emerges through training for the evaluated LLMs, including coupling across depth (Figure 3 ) and across tokens (Figures 4, LABEL:fig:coupling_fixedin, LABEL:fig:coupling_fixedout). Further, we evaluate layer-wise coupling at intermediate training checkpoints of Pythia 6.9B and 12B (Biderman et al., 2023) (Figure 1b), and observe that coupling is generally low at initialization and increases persistently throughout training. Moreover, there is a clear sense of locality in the strength of coupling which is visually displayed in Figure (1c-e)
The gradual growth of coupling observed in Figure (1b) parallels the logarithmic increase in accuracy during training for Pythia 12B and Pythia 6.9B (Biderman et al., 2023). This highlights the relationship between coupling and performance, since both properties emerge at similar training iterations and rates.
@@ 5.4 Correlation with Generalization
For each LLM, we measure the average coupling across depth(which we define to be negative mis-coupling) across prompts in the 6 evaluation datasets (Section 4.2), where for each prompt, $m(A_{ll^{\prime},K}^{n})$ is averaged over layers $l,l^{\prime}\in\{1,\ldots,L\}$. We plot the coupling values against the benchmark scores across several LLMs (Figure 1). Our results reveal a positive correlation between coupling and performance benchmark scores, and is more significant than the relationship between other significant model hyperparameters (Figure LABEL:fig:correlation_hyperparams). This observation suggests a compelling relationship between stronger coupling of singular vectors of Jacobians $J_{l}^{t_{1}t_{2}}$ and improved generalization.
The results for ViTs demonstrate that stochastic depth encourages coupling during training (Figure LABEL:figure:coupling_vitb) and that coupling correlates with accuracy when fixing SD rate (Figure LABEL:figure:coupling_vita). This finding suggests that coupling may provide new insight into stochastic depth’s underlying mechanism, and that developing training methods to amplify coupling across transformer blocks could provide additional regularization and improve performance.
We hypothesize that simple trajectories may lead to better generalization. This agrees with many generalization bounds in machine learning (Arora et al., 2018), which suggest that models with lower complexity tend towards better generalization. Additionally, prior works (Novak et al., 2018) demonstrate that the Frobenius norm of input-output Jacobians is related to generalization, providing evidence that coupling — a structural property derived from Jacobians — may also correlate with generalization.
5.5 Regularity in Hidden Trajectories
We identify a considerable degree of linearity in the hidden trajectories of all featured LLMs. The LSS of the trajectories has an average value $4.25$ in trained models, while taking average values of $6.54$ at initialization, and is supported by the low variation across benchmark prompts (Figure LABEL:figs:lss-expo). Linearity increases with training at varying depths of Pythia 12B (Figure LABEL:figs:lss-training). Linear and expansive behavior of the representations is demonstrated in Llama-3 70B, MPT 30B, and NeoX 20B through low dimensional projections of embedding trajectories (Figure LABEL:figs:pca).
Among the LLMs considered, the vast majority of hidden trajectories exhibit exponential growth that emerges with training (Figure LABEL:fig:expo-all b). Many models exhibit a low coefficient of variation across prompts, showing the robustness of this property across a variety of tasks. In contrast, measurements at initialization show equally (rather than exponentially) distanced trajectories, as reflected by the low coefficients of variation in the norms of their layer-wise differences (Figure 23). Under certain assumptions, exponential spacing is motivated in Section 6.1.

### v4

@@ Abstract
Large Language Models (LLMs) have made significant strides in natural language processing, and a precise understanding of the internal mechanisms driving their success is essential. In this work, we analyze the trajectories of token embeddings as they pass through transformer blocks, linearizing the system along these trajectories through their Jacobian matrices. By examining the relationships between these block Jacobians, we uncover the phenomenon of transformer block coupling in a multitude of LLMs, characterized by the coupling of their top singular vectors across tokens and depth. Our findings reveal that coupling positively correlates with model performance, and that this relationship is stronger than with other hyperparameters such as parameter count, model depth, and embedding dimension. We further investigate the emergence of these properties through training, observing the progressive development of coupling, as well as increased linearity and layer-wise exponential growth in the token trajectories. Additionally, experiments with ViTs further validate emergence of coupling and its correlation between coupling and generalization, complementing our findings in LLMs. Collectively, these insights provide a novel perspective on token interactions in transformers and open directions for studying and improving training and generalization.
@@ 1 Introduction
Transformers (Vaswani et al., 2017) can be represented as discrete, nonlinear, coupled dynamical systems, operating in high dimensions (Greff et al., 2016; Papyan et al., 2017; Haber & Ruthotto, 2017; Ee, 2017; Ebski et al., 2018; Chen et al., 2018; Bai et al., 2019; Rothauge et al., 2019; Gai & Zhang, 2021; Li & Papyan, 2023). Viewing the skip connections as enabling a discrete time step, we represent the hidden representations as dynamically evolving through the layers of the network. The term nonlinear refers to the nonlinear transformations introduced by activation functions, and coupled refers to the interdependent token trajectories that interact through the MLP and self-attention layers.
In our work, we investigate whether there are identifiable structural characteristics across 30+ pretrained LLMs, analyze their emergence with training, and examine their relationship with generalization performance. During inference, as token embeddings pass through the network, we analyze the their layer-wise trajectories and linearize the effect of transformer blocks on the token embeddings across model depth to study their relationships. To this end, we compute the Jacobians of distinct connections between layers and tokens, perform singular value decompositions (SVDs), and compare the resulting singular vectors. This approach measures the degree of coupling between singular vectors to capture the operational similarity of blocks as they act on tokens. This perspective raises several key questions:
@@ 1.1 Contributions
We investigate the motivating questions across an extensive collection of openly-available LLMs, most having over 1B parameters, trained by over 8 independent organizations, with varying training methods and data (Appendix A.1). Through our experiments, we identify consistent properties that characterize the transformer block coupling phenomenon.
1. Coupling. The singular vectors of the Jacobians of the transformer blocks are coupled across depth (Figures 4, LABEL:fig:paris_all_8to16_appendix) and tokens (Figures 5, LABEL:fig:coupling_fixedin, LABEL:fig:coupling_fixedout, LABEL:fig:coupling_inoutsame_appendix, LABEL:fig:coupling_fixed_input_appendix, LABEL:fig:coupling_fixed_output_appendix) in several open source LLMs (Table (1)). Furthermore, coupling across Jacobians emerges with training (Figures LABEL:fig:pythia-check, 4, 5, LABEL:fig:vit_coupling_v_epoch, LABEL:fig:coupling_fixedin, LABEL:fig:coupling_fixedout), and the coupling strength becomes more pronounced between adjacent layers with training, indicating a layer-wise locality in the interactions (Figures LABEL:fig:coupling-mega-plot(c-e)).
2. Generalization. The strength of coupling is correlated with benchmark performance on the HuggingFace Open LLM Leaderboard (Beeching et al., 2023) (Figures LABEL:fig:llm-correlation, LABEL:fig:alignment_vary_k). Additionally, coupling is more strongly correlated with generalization than parameter budget, model depth, and token embedding dimension (Figure LABEL:fig:correlation_hyperparams). Experiments with ViTs support the relationship observed in LLMs (Figure LABEL:fig:vit_acc_v_coupling).
3. Regularity. Linearity in hidden trajectories emerges with training (Figures LABEL:figs:lss-training, LABEL:figs:pca, LABEL:fig:lss-all, 14, 19, 20), aligning with behaviour previously observed in ResNets Li & Papyan (2023). Exponential growth occurs in contiguous token representations as a function of depth (Figures LABEL:fig:expo-all, 18, 19, LABEL:figs:expo-training), starkly contrasting linear growth previously observed.
We provide a new perspective on token embedding interactions by examining layers of transformers through their Jacobian matrices. Our results display the effects of training on transformer blocks, and opens potential avenues for promoting generalization in similar architectures.
@@ 2 Background on Large Language Models
LLMs may be described as a deep composition of functions that iteratively transform token embeddings. In the input layer, $l=0$, the text prompts are tokenized and combined with the positional encodings to create an initial high-dimensional embedding, denoted by $x_{i}^{0}\in\mathbb{R}^{d_{\text{model}}}$ for the $i^{\text{th}}$ token. When these embeddings are stacked, they form a matrix:
@@ 3.1 Coupling of Singular Vectors of Jacobians
$J_{t_{1}t_{2}}^{l}=U_{l}S_{l}V_{l}^{\top}$ and $J_{t_{1}^{\prime}t_{2}^{\prime}}^{l^{\prime}}=U_{l^{\prime}}S_{l^{\prime}}V_{l^{\prime}}^{\top},$
@@
$A_{t_{1}t_{2}t_{1}^{\prime}t_{2}^{\prime},K}^{ll^{\prime}}$ :-$\coloneq U_{l^{\prime},K}^{\top}J_{t_{1}t_{2}}^{l}V_{l^{\prime},K}:$
$=U_{l^{\prime},K}^{\top}U_{l}S_{l}V_{l}^{\top}V_{l^{\prime},K},$
for $l,l^{\prime}\in\{1,\ldots,L\}$ and $t_{1},t_{2},t_{1}^{\prime},t_{2}^{\prime}\in\{1,\ldots,n\}$, where $U_{l^{\prime},K}$ and $V_{l^{\prime},K}$ are the matrices of top $K$ singular vectors. If the singular vectors of distinct Jacobians are strongly aligned, then the coupling matrix $A$ should be strongly diagonal and approximately equal to the matrix, $S_{l,K}$, of the top $K$ singular values of $J^{l}$. To quantify the top $K$ mis-coupling, we define
$m_{K}(J_{t_{1}t_{2}}^{l},J_{t_{1}^{\prime}t_{2}^{\prime}}^{l^{\prime}})=\frac{\lVert A_{t_{1}t_{2}t_{1}^{\prime}t_{2}^{\prime},K}^{ll^{\prime}}-S_{l,K}\rVert_{F}}{\lVert{\bm{s}}_{l,K}\rVert_{p}},$ (9)
where ${\bm{s}}_{l,K}$ is the vector of the top $K$ singular values of $J^{l}$, $\lVert\cdot\rVert_{F}$ denotes the Frobenius norm and $\lVert\cdot\rVert_{p}$ the vector $p$-norm. Typically, we take $p=1$, which corresponds to normalizing by the sum of the top $K$ singular values of $J^{l}$. We then define the top $K$ coupling to be $c_{K}=1-m_{K}$.
We also generalize this analysis to the alignment between left and right singular vectors of different Jacobians by considering the matrix
$B_{t_{1}t_{2}t_{1}^{\prime}t_{2}^{\prime},K}^{ll^{\prime}}$ :-$\coloneq V_{l^{\prime},K}^{\top}J_{t_{1}t_{2}}^{l}U_{l^{\prime},K}.:$
@@ 3.3 Layer-Wise Exponential Growth
This measurement is motivated by the discussion in Section 6.1, the parametrization in Appendix A.5, and empirical evidence from Figure LABEL:figs:norm-expo-traininga. It also serves as a method to assess the validity of the linearization presented in Equation 14.
@@ 3.4 Vision Transformer Training
For further investigation of coupling in transformers, we consider Vision Transformers (ViTs) following DEiT (Touvron et al., 2021). We train 3 ViTs for classification on CIFAR10 (Krizhevsky, 2009) with varied stochastic depth rate for embedding size 192, depth 24, and 3 attention heads. Please see Appendix A.7 for further details.
Figure 4: Transformer Block Coupling across Depth. The figure shows Jacobian coupling across transformer blocks 9 to 16, using the prompt ”What is the capital of France? The capital is” to trace the final token’s trajectory. In trained models (bottom row), the diagonal pattern with minimal off-diagonal values indicates alignment of Jacobians, where top singular vectors of $J^{l^{\prime}}$ diagonalize $J^{l}$. Untrained models do not exhibit such coupling. Further details are in the Appendix A.8 (Figure LABEL:fig:paris_all_8to16_appendix). Best viewed in color.
@@ 4.1 Suite of Large Language Models
Our study evaluates a comprehensive collection of LLMs (see Appendix A.1) that were independently trained by various individuals and organizations. These models, provided through HuggingFace (Wolf et al., 2020), vary in terms of parameter budgets, number of layers, hidden dimensions, and training tokens. Moreover, we analyze the dynamics of each measurement throughout training by deploying the Pythia Scaling Suite (Biderman et al., 2023). A summary of the models under consideration is presented in Table 1 of Appendix A.1 and further details in Appendix A.6.
@@ 4.2 Prompt Data
We evaluate these LLMs using prompts of varying length, ambiguity, and context, sourced from the test set of ARC (Clark et al., 2018), GSM8K (Cobbe et al., 2021), HellaSwag (Zellers et al., 2019), MMLU (Hendrycks et al., 2021), Truthful QA (Lin et al., 2022), and Winogrande (Sakaguchi et al., 2019). These set the performance benchmarks on the HuggingFace Open LLM Leaderboard (Beeching et al., 2023) and provide a representative evaluation of performance on many language tasks.
@@ 5.1 Coupling of Jacobians Across Depth
In trained LLMs, we observe coupling of the top singular vectors of the Jacobians across depth (Figure 4 bottom row), evident in the low non-diagonal values with a visible diagonal present in the matrix subplots. This is consistently observed across various LLMs considered. On the other hand, in untrained models (Figure 4 top row), there is no coupling of Jacobians across different depths. There is coupling along the diagonal, however, because each Jacobian is trivially diagonalized by its own singular vectors. This, in addition to Figure LABEL:fig:pythia-check, suggests that coupling across depth emerges through training.
@@ 5.2 Coupling of Jacobians Across Tokens
We analyze the coupling of singular vectors of Jacobians across tokens. For input and output tokens that are the same ($J_{l}^{tt}$ and $J_{l^{\prime}}^{t^{\prime}t^{\prime}}$, Figure 5), we observe strong coupling, indicating that a token’s interactions along its trajectory are coupled with others. For context tokens, coupling is examined by fixing the input token ($J_{l}^{t_{1}t_{2}}$ and $J_{l^{\prime}}^{t_{1}t_{2}^{\prime}}$, Figure LABEL:fig:coupling_fixedin) or the output token ($J_{l}^{t_{1}t_{2}}$ and $J_{l^{\prime}}^{t_{1}^{\prime}t_{2}}$, Figure LABEL:fig:coupling_fixedout). While context coupling exists, its strength varies across token pairs. Untrained models show no such coupling with randomly fixed layers $l,l^{\prime}$ in general.
Figure 5: Transformer Block Coupling across Tokens (Self Coupling). The figure shows Jacobian coupling for the same input and output token across tokens, visualized using the absolute values of $A_{ll^{\prime}}^{ttt^{\prime}t^{\prime}}$ (with fixed layers $l,l^{\prime}$). In trained models (bottom row), the strong diagonal and small off-diagonal values indicate coupling, while no such coupling is present at initialization (top row). Additional plots are in Appendix A.8 (Figure LABEL:fig:coupling_inoutsame_appendix).
@@ 5.3 Emergence of Coupling with Training
Coupling emerges through training for the evaluated LLMs, including coupling across depth (Figure 4) and across tokens (Figures 5, LABEL:fig:coupling_fixedin, LABEL:fig:coupling_fixedout). Further, we evaluate layer-wise coupling at intermediate training checkpoints of Pythia 6.9B and 12B (Biderman et al., 2023) Figure LABEL:fig:pythia-check, and observe that coupling is generally low at initialization and increases persistently throughout training. Moreover, there is a clear sense of locality in the strength of coupling which is visually displayed in Figure LABEL:fig:coupling-mega-plotc-e)
The gradual growth of coupling observed in Figure LABEL:fig:pythia-check parallels the logarithmic increase in accuracy during training for Pythia 12B and Pythia 6.9B (Biderman et al., 2023). This highlights the relationship between coupling and performance, since both properties emerge at similar training iterations and rates.
5.4 Coupling in Vision Transformers
To further investigate coupling, we train several ViTs on CIFAR-10. Our experiments confirm that coupling emerges with training (Figure LABEL:fig:vit_coupling_v_epoch) and exhibits a strong positive correlation with test accuracy (Figure LABEL:fig:vit_acc_v_coupling). These findings suggest that coupling is a general phenomenon in transformers and reinforces our observations in LLMs. Notably, we measure coupling in ViTs on train data while its correlation with accuracy persists on test data, highlighting coupling as a structural property of the model related to generalization.
@@ 5.5 Correlation with Generalization
For each LLM, we measure the average coupling across depth for prompts in the 6 evaluation datasets (Section 4.2), where for each prompt, $c_{K}(J_{n}^{l},J_{n}^{l^{\prime}})$ is averaged over layers $l,l^{\prime}\in\{1,\ldots,L\}$. We then plot the coupling values against the benchmark scores across several LLMs (Figure LABEL:fig:llm-correlation). Our results reveal a positive correlation between coupling and performance benchmark scores, exceeding the significance of correlations with other model hyperparameters (Figure LABEL:fig:correlation_hyperparams). This observation suggests a compelling relationship between stronger coupling of singular vectors of Jacobians $J_{l}^{t_{1}t_{2}}$ and improved generalization.
We hypothesize that simple trajectories may lead to better generalization. This idea aligns with with several generalization bounds in machine learning (Arora et al., 2018), which suggest that models with lower complexity tend towards better generalization. Furthermore, previous work (Novak et al., 2018) demonstrated that the Frobenius norm of input-output Jacobians is related to generalization, suggesting that coupling, a structural property derived from block Jacobians, may also be related to generalization.
5.6 Regularity in Embedding Trajectories
By analyzing token trajectories as representations propagate through the layers, we identify several structural properties. Notably, we observe a considerable degree of linearity in the hidden trajectories across all examined LLMs. The LSS of the trajectories has an average value $4.25$ in trained models, compared to $6.54$ at initialization, and is supported by the low variation across benchmark prompts (Figure LABEL:figs:lss-expo). Linearity increases with training at varying depths of Pythia 12B (Figure LABEL:figs:lss-training). This linear and expansive behavior of the representations is further illustrated in Llama-3 70B, MPT 30B, and NeoX 20B through low dimensional projections of embedding trajectories (Figure LABEL:figs:pca).
Among the LLMs studied, most hidden trajectories exhibit exponential growth that emerges with training (Figure LABEL:fig:expo-all). The low coefficient of variation across prompts indicates the robustness of this property across diverse tasks. In contrast, measurements at initialization show equally (rather than exponentially) distanced trajectories, as reflected by the low coefficients of variation in the norms of their layer-wise differences (Figure 18). Under certain assumptions, exponential spacing is motivated through coupling, as discussed in in Section 6.1 and Appendix A.5.